OpenStudy (anonymous):

What is the equation of the line that is parallel to y = –4x + 3 and has a y-intercept of -1/3?

OpenStudy (anonymous):



OpenStudy (amistre64):

parallel lines have the same equations ... they differ only by a constant y = mx+b is parallel to y=mx+k

OpenStudy (anonymous):

Well if it parallel then it has the same slope. y=mx+b m=slope b=y-intercept So your slope has to be -4. And your y-intercept must be -1/3 So just plug in the stuff into this equation y=mx+b